# Water vapor emission at the warm cavity wall of the HD 100546 disk as revealed by ALMA

### v1

@@ 1 Introduction
In this work we present the first spatially resolved ALMA observations of water vapor in a disk hosting a wide dust cavity. The source is described in Sect. 2, while observations used to perform the analysis, self-calibration and image reconstruction are outlined in Sect. 3. We discuss the main results in Sect 4, and summarize the main conclusions in Sect. 5.
2 The source
@@
L. M. Pirovano et al. (2022) modeled the Herschel observations of HD 100546 using DALI (S. Bruderer et al., 2012; S. Bruderer, 2013) and constrained the water abundance in the three reservoirs discussed in Sect. 1. Their fiducial model predicts low water abundances in the thermally desorbed component () and in the hot reservoir formed in the gas phase (), while the photodesorbed component at intermediate layers outside the dust ring at au needs to be relatively abundant () to reproduce the observations.
3 Observations
@@ 3.2 Imaging
3.3 Moment maps, radial profile and spectrum
@@
We show the radial profile obtained from the cube imaged with a robust parameter of 0.5 in favor of a higher spatial resolution to highlight possible radial features. The inner peak of the water emission coincides with the inner edge of the dust cavity wall, as visible from the comparison of the solid blue and dashed black profiles in Fig. 2. The outer azimuthally asymmetric emission observed in the integrated intensity map in Fig. 1 is also visible in the radial profile in Fig. 2, which shows an outer emission bump peaking at au. We will further discuss the emission morphology of water vapor in Sect. 4.
@@
4 Discussion
The presented observations show the first spatially resolved detection of the main water vapor isotopologue with ALMA in a transition disk. As anticipated in Sect. 3.3, the water emission at 183.31 GHz shows an inner bright peak at au. In the following sections we discuss this result in light of the previous dust and line observations of the disk around HD 100546.
4.1 Water vapor at the warm cavity wall
@@
4.2 Comparison with Herschel observations
@@
The hot water lines with high K) are expected to trace the hot water reservoir produced in the gas phase in the upper layers and inside the dust cavity (see Sect. 1), where photodissociation limits the water abundance. Water self-shielding has been shown to mitigate this effect when the water column density is high ( cm-2, T. Bethell & E. Bergin 2009; A. D. Bosman et al. 2022; J. K. Calahan et al. 2022). However, HD 100546 is a transition disk with a large dust cavity which is also depleted in gas. The right panel of Fig 6 in Appendix C represents the radial profile of the water column density from the thermochemical model by M. Leemker et al. (2024), which shows that the water column density in the disk cavity is expected to be low enough to allow efficient photodissociation. We highlight that this model includes photodissociation only in the vertical direction, and not in the radial one. The non-detection of the hot Herschel lines with K which trace the hot water reservoir in the upper layers and inside the cavity can be thus explained by the efficient photodissociation (L. M. Pirovano et al., 2022).
@@
4.3 Water vapor emission and the molecular complexity of HD 100546
@@ Appendix B Water channel maps
Figure 5 shows the channel maps of the detected water line in the HD 100546 disk. The cube was imaged with natural weighting and a channel width of km s(see Sect. 3.2 for more details on the imaging procedure).

### v2

@@ I Introduction
In this work we present the first spatially resolved ALMA observations of water vapor in a disk hosting a wide dust cavity. The source is described in Sect. II, while observations used to perform the analysis, self-calibration and image reconstruction are outlined in Sect. III. We discuss the main results in Sect IV, and summarize the main conclusions in Sect. V.
II The source
@@
L. M. Pirovano et al. (2022) modeled the Herschel observations of HD 100546 using DALI (S. Bruderer et al., 2012; S. Bruderer, 2013) and constrained the water abundance in the three reservoirs discussed in Sect. I. Their fiducial model predicts low water abundances in the thermally desorbed component () and in the hot reservoir formed in the gas phase (), while the photodesorbed component at intermediate layers outside the dust ring at au needs to be relatively abundant () to reproduce the observations.
III Observations
@@ III.2 Imaging
III.3 Moment maps, radial profile and spectrum
@@
We show the radial profile obtained from the cube imaged with a robust parameter of 0.5 in favor of a higher spatial resolution to highlight possible radial features. The inner peak of the water emission coincides with the inner edge of the dust cavity wall, as visible from the comparison of the solid blue and dashed black profiles in Fig. 2. The outer azimuthally asymmetric emission observed in the integrated intensity map in Fig. 1 is also visible in the radial profile in Fig. 2, which shows an outer emission bump peaking at au. We will further discuss the emission morphology of water vapor in Sect. IV.
@@
IV Discussion
The presented observations show the first spatially resolved detection of the main water vapor isotopologue with ALMA in a transition disk. As anticipated in Sect. III.3, the water emission at 183.31 GHz shows an inner bright peak at au. In the following sections we discuss this result in light of the previous dust and line observations of the disk around HD 100546.
IV.1 Water vapor at the warm cavity wall
@@
IV.2 Comparison with Herschel observations
@@
The hot water lines with high K) are expected to trace the hot water reservoir produced in the gas phase in the upper layers and inside the dust cavity (see Sect. I), where photodissociation limits the water abundance. Water self-shielding has been shown to mitigate this effect when the water column density is high ( cm-2, T. Bethell & E. Bergin 2009; A. D. Bosman et al. 2022; J. K. Calahan et al. 2022). However, HD 100546 is a transition disk with a large dust cavity which is also depleted in gas. The right panel of Fig 6 in Appendix C represents the radial profile of the water column density from the thermochemical model by M. Leemker et al. (2024), which shows that the water column density in the disk cavity is expected to be low enough to allow efficient photodissociation. We highlight that this model includes photodissociation only in the vertical direction, and not in the radial one. The non-detection of the hot Herschel lines with K which trace the hot water reservoir in the upper layers and inside the cavity can be thus explained by the efficient photodissociation (L. M. Pirovano et al., 2022).
@@
IV.3 Water vapor emission and the molecular complexity of HD 100546
@@ Appendix B Water channel maps
Figure 5 shows the channel maps of the detected water line in the HD 100546 disk. The cube was imaged with natural weighting and a channel width of km s(see Sect. III.2 for more details on the imaging procedure).